(x-3)(x^6-2x^3+1)=0
Answer by stanbon(75887)

(x-3)(x^6-2x^3+1)=0
(x-3)(x^3-1)^2 = 0
(x-3)(x-1)(x^2+x+1) = 0
------------
Roots of x^2+x+1:
x = [-1 +- sqrt(1-4*1*1)]/2
x = [-1 +- isqrt(3)]/2
====================
Ans: x=3 or x=1 or x = [-1 +- isqrt(3)]/2
====================
Cheers,
Stan H.